A culture of Bacteria obeys the law of uninhibited growth. if 500 bacteria are present initially and there are 800 after 1 hour, how many will be present in the culture after 5hours? How long is it until there are 20,000 bacterias?
Answer by Boreal(15235)

A=Aob^t
800=500*b^1
b=8/5=1.6
A5=500(1.6)^5=5242.88=5243
20000=500*1.6^t
40=1.6^t
log 40=t log 1.6
log 40/log 1.6=t=7.85 hours